# references.bib

### v1

@@ II RELATED WORK
Learning compliance from Reinforcement Learning (RL). RL can effectively learn compliance controllers by exploring force-motion variations [portela2024learning, kalakrishnan2011learning, beltran2020learning, beltran2020variable, noseworthy2024forge, chang2022impedance]. However, these policies need to be retrained for any scene variations. Moreover, most of existing work uses fixed-parameter low-level compliance controllers like impedance [lee2019making] or admittance [kohler2024symmetric] controllers, which lack robustness against disturbances.
Learning compliance from human Human stiffness during a manipulation task can be estimated from sufficient repetitions of the same motion [deng2016learning, duan2018learning, duan2019sequential, wang2020framework, yamane2023soft, zeng2021generalization]. The stiffness is considered to be either proportional to the force covariance [duan2018learning, duan2019sequential] or inversely proportional to the position covariance [calinon2010learning, kronander2012online]. These methods typically do not work for visuomotor policy learning since every demonstration is different. [grioli2011non] proposed a stiffness estimator that works with one single demonstration. However, it requires perfect knowledge of human mass and damping, which is hard to obtain. Prior work also attempted to give the demonstrator the ability to specify stiffness [kronander2013learning, kamijo2024learning] without spatial variation.
@@ III-A Preliminaries: Modeling
Manipulation Modeling Next we model how a manipulation system interacts with a robot, treating the robot as a black box. We make the following assumption:
@@ IV-B Task II: Vase Wiping
Compared with the [Compliant Policy], [ACP] maintains accurate tracking of the desired motion for being stiff in the motion directions, as shown in Fig. 8, top row. The wiping motion of the [Compliant Policy] deviates from the position target because it is sensitive to the friction from the vase. As a result, the sliding motion is insufficient to make a high quality wipe. The comparison of wiping results is shown in Tab. II, which excludes the [Stiff Policy] since it broke the robot tool frequently.
@@ V CONCLUSIONS
In this work, we show that Adaptive Compliance Policy is an effective visuomotor policy to do compliant manipulation. Extensive result in real-world shows that our approach is able to extract useful compliance from human demonstration in a variety of manipulation tasks, and thereby significantly improve the success rate of two contact-rich manipulation tasks.

### v2

@@ II RELATED WORK
Learning compliance from Reinforcement Learning (RL). RL can learn compliance controllers by exploring force-motion variations [portela2024learning, kalakrishnan2011learning, beltran2020learning, beltran2020variable, noseworthy2024forge, chang2022impedance, martin2019variable]. However, these policies need to be retrained for any scene variations. Moreover, most of existing work uses fixed-parameter low-level compliance controllers like impedance [lee2019making] or admittance [kohler2024symmetric] controllers, which lack robustness against disturbances.
Learning compliance from human Human stiffness during manipulation can be estimated from sufficiently repeated same motions [deng2016learning, duan2018learning, duan2019sequential, wang2020framework, yamane2023soft, zeng2021generalization]. The stiffness is either proportional to the force covariance [duan2018learning, duan2019sequential] or inversely proportional to the position covariance [calinon2010learning, kronander2012online]. These methods do not work for visuomotor policy learning since every demonstration is different. [grioli2011non] proposed a stiffness estimator that works with a single demonstration. However, it requires perfect knowledge of human mass and damping, which is impractical. Prior work also attempted to give the demonstrator the ability to specify uniform stiffness [kronander2013learning, kamijo2024learning]. A concurrent work [liu2024forcemimic] computes compliance profiles based on heuristics from manually divided stages of tasks.
@@ III-A Preliminaries: Modeling
Manipulation Modeling. Next we model how a manipulation system interacts with a robot, treating the robot as a black box. We make the following assumption:
@@ IV-B Task II: Vase Wiping
Compared with the [Compliant Policy], [ACP] maintains accurate tracking of the desired motion, as shown in Fig. 8, top row. The wiping motion of the [Compliant Policy] deviates from the position target because it is sensitive to the friction from the vase. As a result, the sliding motion is insufficient to make a high quality wipe. The comparison of wiping results is shown in Tab. II, which excludes the [Stiff Policy] since it broke the robot tool frequently.
@@
We also compare with a policy that predicts force and position, which is less robust without a suitable compliance profile. See our website for details.
@@ V CONCLUSIONS
In this work, we show that Adaptive Compliance Policy is an effective visuomotor policy for compliant manipulation. Extensive real-world results show that our approach is able to extract useful compliance from human demonstration, and thereby significantly improve the success rate of two contact-rich manipulation tasks.